style=”color:black”>Journalist Report Feb 25th
Crew 206 – 02/25/2019
Author : Auzou Benjamin, Journalist
Title : "Good Morning Mars !"
0847 : The airlock opens, and the crew 206 is in front of a whole new world to explore : Mars.
The marswalkers of the day, Jérémy, Cerise, Aurélien and Norbert leaved the Hab to explore the surroundings of the station and to find a proper place for the LOAC, an experiment we will deploy tomorrow.
But above all that was for three of them their first steps on the Red Planet. I think they’re better than me to describe the feelings they had when they touched the martian ground for the first time.
"This first EVA was a really special and exciting moment. Everything went good, and I just want to go again on EVA !" – Norbert
"I really felt like it was my first step on Mars, and it is an
amazing feeling." – Cerise
"This was an exiting moment. The team was super pro and it made the
experience amazing" – Aurélien
Before that, the morning was really busy. We woke up very early in order to begin the EVA at 0830 so that our 4 marswalkers avoid the mud. We beginned the day with a sport session and then we prepared our first martian pancakes to bear the morning’s program. I was among the
three members of the crew that followed the intense session, seven different exercices during fourteen minutes, it was hard, especially the push-ups. However it is important to stay active, and our daily performances are monitored by our Commander.
During the EVA, I was assigned by Norbert to take care of the Green Hab. That place is amazingly calm and peaceful. I watered all the plants : tomatoes, basil, cress, radishes; they need a lot of water, 8 liters this morning for the whole Green-Hab but they worth it !
Our Astronomer Aurélien started his first observations of the sun. This is primordial for us on Mars, without a decent atmosphere, the solar radiations are a serious danger !
Today was also the beginning for three experiments : Norbert’s radishes are now planted, Cerise took some samples of the drinking water and the whole crew spend fifteen minutes on our Virtual Reality experiment under a heart monitoring : half of us watching entertaining
videos and the other half as witness subjects.
The Hab is now our home, and we continue to make it better and ours. The main improvement of the day is a live plan of the sol displayed in the hab, detailling in real time the tasks for each of us, but also our data consumption.
After this day of hard work, we all need some time for us to rest and decompress. That’s why we chose to watch a movie together tonight : Le Grand Bazar by Les Charlots, a French comedy. That’s my suggestion, I’ll tell you tomorrow if everybody liked it.
style=”color:black”>Operations Report Feb 25th
Crew 206 Operations Report 25 – FEB – 2019
Name of person filing report: Gaspard Thieulin
Non-nominal systems: Generator
Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.
Generator (hours run): Turned off at 0830 (24th Feb), Turned on at
1600 (24th Feb)
Solar— SOC % – Turned off Gen (24th Feb) 76%, Turned on Gen (24th Feb) 82%
Diesel Reading – ~ 75%
Propane Reading – ~ 25%
Ethanol Free Gasoline – 10 gallons
Water (Auxillary tank) – Not in use
Water (Static tank) – around 95%, ~ 522 gallons
Auxillary to Static tank transfer – No
Gallons transferred: Not Applicable
Water in GreenHab – ~ 293,7 gallons
Water (loft) – 12 marked level
Static to Loft Pump used – yes
Water Meter: 01415594 units
Toilet tank emptied: No
Deimos rover used: No, still not functional
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable
Opportunity rover used: Used
Hours: 51.5 hours
Beginning charge: 100%
Ending charge: 74%
Currently charging: Yes
Curiosity rover used: Used
Hours: 75.8 hours
Beginning charge: 100%
Ending charge: 75.5%
Currently charging: Yes
Spirit rover used: Not Used
Hours: 78.3 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: NO
Notes on rovers: Opportunity’s socket is having issues that even the reset trick doesn’t solve. Because it’s currently more viable than Spirit, I unplugged the latter to plug in Opportunity. Charge level was 76% for Opportunity in consequence.
ATV’s Used: N/A
Reason for use: Not used
Oil Added? N/A
ATV Fuel Used: N/A
# Hours the ATVs were Used today: None
Notes on ATVs: None
HabCar used and why, where? Not used
Crew Car used and why, where? Not used
General notes and comments: We are experiencing a lot of static charges in the Hab. We don’t exactly know what is the reason behind that; but it might be a wire not properly connected to the ground.
Summary of internet: Lost Internet access twice in the day. Resetting the Nighthawk router did the trick without any issue.
Summary of suits and radios: Visor fogging on EVA#1 – check suit air supply
Summary of Hab operations: Nothing to report
Summary of Green Hab operations: Nothing to report
Summary of Science Dome operations: Nothing to report
Summary of RAM operations: Pre-built the LOAC and meteo station experiments. We ask to use one clamp to fix the LOAC for the duration of the experiment.
Summary of any observatory issues: Nothing to report.
Summary of health and safety issues: Nothing to report
Questions, concerns and requests to Mission Support: Propane tank still needs to be filled rapidly, how do we proceed ?
style=”color:black”>Green Hab Report Feb 25th
Crew 206 Green Hab Report – 25-FEB-2019
GreenHab Officer: Norbert Pouzin
Environmental control: Heating and Cooling w/ ambient air (0 hrs)
Shade Cloth (40%)
Hours of supplemental light: 5 Hours
Daily water usage for crops: 6.3 gallons
Water in Blue Tank: ~ 293.7 gallons
Crops watered at 0900 hours (2.1 gallons) and 1830 hours (4.2 gallons):
TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
1) 09.00 AM 24.4° C / 21 %
2) 11.30 AM 29.4° C / 0 %
3) 15:00 PM 23.3° C / 0 %
4) 18:00 PM 19.6° C / 27 %
Changes to crops: I planted 8 pots of radish, as part of my plant experiment. I put them in a shelf and moved some tomatoes plant to the other shelf.
Narrative: Yesterday night, the blue tank was refill so we have a full water tank for the greenhab.
In the morning I changed the yellow fly papers.
This afternoon, I started the installation of my plant experiment. I planted 8 pots of radish (Raphanus Sativus), following my protocol. I tried the UV lamp for the experiment, but it doesn’t work. We tried to fix it at the RAM, but it seems that the fact that the frequency used for electricity is deferent from Europe makes it impossible to work. Thus the base we have needs to be changed. I think our UV bulb works properly.
As yesterday, I opened the door from 1200 to 1700 to cool down the greenhab.
7.0g of basil and 1.1g of parsley were harvested for soup.
Harvest: 7.0 g basil and 1.1 g of parsley.
Support/supplies needed: A UV lamp, or a base for the UV bulb we have.
style=”color:black”>Astronomy Report Feb 25th
Astronomy Report, Sol 1
Name: Aurélien Mure
Solar Features Observed: None, tiny clouds let me show just the shape of the sun
Problems encountered : Problems with the automated sun tracker, solved.
EVA Report Feb 25th
Crew 206 EVA Report 25 Feb 2019
Author : Jeremy Auclair
Purpose of EVA : Training new crew members and finding a spot to install the LOAC and the weather station
Start time : 08:42 am
End time : 10:11 am
Narrative : After the 5 minute depressurization, we checked the water levels and the rovers’ batteries. At 08:52 am, We took Opportunity and Curiosity to go towards the Stream Bed Connector intersection. On the way, we stopped at 08:55 at the spot where we had put the LOAC last year (crew 189), to have a comparison point. We then went to Pooh’s corner at 08:58 to have a small walk (7-8 minutes). At this point my visor had started to get pretty foggy (with backpack 1), but I could still walk around safely. We the took the rovers again to head to the Stream Bed Connector, but with the snow we failed to recognize it. We went a bit to far (almost up to Galileo road). After tunrning around we arrived at 09:15 am. We walked around for about 5 minutes, once again to get used to the suits. We then headed back to the Hab to park the rovers. At 09:30 am we plugged both rovers in and we started
walking East North-East to find a good location to install our instruments. They have to be in a relatively flat place, not to close to any hills or mounts, and close to the road, to facilitate changing the batteries (the LOAC is powered with car batteries). We found this place at 09:40 am (–). After that we continued walking around and talking picture before heading back to the Hab. We arrived in the airlock at 10:04 am. One minute after pressurization, we realised we had forgot the EVA equipment bag in Curiosity, so we depressurized again, went to retreive it, and got in the airlock one final time. Pressurization started at 10:06 am. We got in the Hab at 10:11 am.
Pooh’s corner : 518800E 4250900N
Stream Bed Connector : 519000E 425100N
Participants : Arélien Mure, Cerise Cuny, Norbert Pouzin and Jérémy
Auclair (EVA leader)
Road used : Cow Dung road
Mode of travel : Rovers and walking
Sol Summary Report Feb 25th
Crew 206 Sol Summary February 25th 2019
Summary Title : Foggy morning, but sunny day
Author : Jeremy Auclair
Mission Status : All nominal
Sol Activity Summary : After our change of schedule for this morning’s EVA, we had to get out of bed at 06:30 am. A few minutes after that we started our daily physical training. We managed to start breakfast around 07:10 am. We had a hard time making pretty pancakes with the old pan we have. We managed to start preparing for the EVA at 07:40 am. This first EVA is mainly used to train the new crew members, but we also looked for a appropriate spot to install our two instruments, LOAC and the weather station. But this EVA was very annoying for me: 10 minutes after the beginning my visor was completely fogged, but we managed to finish the EVA (I didn’t drive obviously). Otherwise the EVA went very smoothly. After lunch, around 02:00 pm, everyone went their own way to prepare their experiments. Cerise worked on Aquapad, to test the microbiological quality of our drinking water, Aurélien went to the solar observatory to get used to it, Norbert worked on his plants. After that we all participated to the virtual reality (VR) experiment. Its goal is to study the efficiency of VR as a countermeasure to confinement and isolation.
Look ahead plan : Tomorrow will be a big day, a lot of our main experiments will be installed and performed for the first time, so we are preparing as much as we can today. The crew is slowly getting used to the life in the station. Water consumption monitoring is heading in the right direction, everyone follows the plan correctly.
Anomalies in work : none
Weather : a little cloudy until 2 pm, after which the sky was quite clear.
Crew physical status : small aches from physical training but fine otherwise
EVA : Pooh’s Corner and Stream Bed Connector
Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA Request
Support Requested : a new frying pan, a new sponge for the dishes, a rag and a tissue napkin to prevent using too much paper towels
You must be logged in to post a comment.